5kg weight who mation is constrained by vertical guides is dropped from reit 0.5 m onto a vertically aligned spring whose stiffness is 10 Nm How fast is the weight travelling when the spring has been compressed by 50 mm

1
Expert's answer
2022-02-23T08:33:07-0500

"mgh=\frac{kx^2}2+\frac{mv^2}2,"

"v=\sqrt{2gh-\frac{kx^2}m}=2.2~\frac ms."